YeetesMeates:

how do u colors work here and levels what does the 18 mean on my pf ??

YeetesMeates:

how does all that work

Shayla:

The 18 is your smart score

Laylalyssa:

MxxnLight:

You have to be more active in posts/answering them. Also early congrats on orange. o:

YeetesMeates:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @MxxnLight You have to be more active in posts/answering them. Also early congrats on orange. o: \(\color{#0cbb34}{\text{End of Quote}}\) thanks

itsmehjay:

Ok so listen loser- when u look on ur profile you will see the diff levels there are 0-19 = orange 20-49 = red 50-74 = purple 75-89 = blue 90-100 = green

MxxnLight:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @YeetesMeates \(\color{#0cbb34}{\text{Originally Posted by}}\) @MxxnLight You have to be more active in posts/answering them. Also early congrats on orange. o: \(\color{#0cbb34}{\text{End of Quote}}\) thanks \(\color{#0cbb34}{\text{End of Quote}}\) Anytime. (:

itsmehjay:

great y'all already beat meee dfhyfnmyj e.e

YeetesMeates:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @itsmehjay Ok so listen loser- when u look on ur profile you will see the diff levels there are 0-19 = orange 20-49 = red 50-74 = purple 75-89 = blue 90-100 = green \(\color{#0cbb34}{\text{End of Quote}}\) X'D

YeetesMeates:

how do i get smart is the queestion

itsmehjay:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @YeetesMeates how do i get smart is the queestion \(\color{#0cbb34}{\text{End of Quote}}\) stay in skool? B)

YeetesMeates:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @itsmehjay \(\color{#0cbb34}{\text{Originally Posted by}}\) @YeetesMeates how do i get smart is the queestion \(\color{#0cbb34}{\text{End of Quote}}\) stay in skool? B) \(\color{#0cbb34}{\text{End of Quote}}\) X'D NO POS WOW

Shayla:

by doing these things it'll improve your smart score

YeetesMeates:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Shayla by doing these things it'll improve your smart score \(\color{#0cbb34}{\text{End of Quote}}\) sounds good thanks

Astrid1:

The numbers on besides your profile picture is your smart score. It will go up if you follow the rules and it can go down if you don't follow the rules. It does take time though, so stay patient. The color will change as your smart score goes up. All you have to do is help others responsibly (like not giving direct answers) and if someone were to make a spam post just click the report button and be on your way. You can also stay active in chats and that can raise your ss.

Shayla:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @YeetesMeates \(\color{#0cbb34}{\text{Originally Posted by}}\) @Shayla by doing these things it'll improve your smart score \(\color{#0cbb34}{\text{End of Quote}}\) sounds good thanks \(\color{#0cbb34}{\text{End of Quote}}\) ofc